# Revisiting Multivariate Time Series Forecasting with Missing Values

### v2

@@ Appendix D Full Experiments
Appendix E Extra Experiments
E.1 Forecasting Results Visualization
We present a spatial visualization of forecasting results to demonstrate the effectiveness of CRIB under varying missing rates. Fig. 5 shows the final timestamp in the historical time window and the first forecasting timestamp on the PEMS-BAY dataset. At lower missing rates (20% and 40%), by effectively leveraging inter-variate correlations extracted from the data, CRIB accurately predicts the future values. Even at higher missing rates (60% and 70%), CRIB can maintain stable performance and predict the spatial distribution of the PEMS-BAY datasets. These findings underscore CRIB’s capability to handle incomplete data and produce reliable predictions.
E.2 Unified-Variate Attention Maps Visualization
In Fig. 6, we compare visualizations of directly applying IB on the Transformer with our proposed CRIB. In the first experiment, a transformer model serves as the predictor. The left two figures clearly show that directly applying IB to the model would force the model to focus on a few specific values (straight line attention), thereby neglecting global information. In contrast, the right figures reveal that CRIB can not only capture the original intra-variate temporal correlations in one attention head but also effectively uncovers cross-variate correlations in another, rather than relying solely on raw correlations. As a result, the final forecasting performance is improved remarkably by our unified-variate attention mechanism and consistency regularization scheme.
E.3 Experiments on Various Missing Patterns

### v3

@@ Appendix H Full Experiments
Table 16: Performance comparison on six additional datasets with varying missing rates. The best results are highlighted in Bold.
Data
Metric SegRNN WPMixer iTransformer PatchTST DLinear PAttn CSDI NeuralCDE ImputeFormer TimesNet Ours
ETTh2 MAE@0% 0.096 0.095 0.098 0.096 0.098 0.096 0.113 0.197 0.101 0.097 0.094
MSE@0% 0.024 0.024 0.025 0.024 0.025 0.025 0.033 0.074 0.026 0.025 0.023
MAE@20% 0.115 0.116 0.115 0.109 0.142 0.116 0.255 0.244 0.122 0.121 0.107
MSE@20% 0.031 0.030 0.033 0.030 0.040 0.033 0.192 0.105 0.034 0.033 0.028
MAE@40% 0.148 0.147 0.298 0.215 0.181 0.289 0.404 0.276 0.140 0.153 0.131
MSE@40% 0.047 0.045 0.180 0.125 0.065 0.173 0.399 0.134 0.049 0.048 0.039
MAE@60% 0.183 0.202 0.270 0.216 0.253 0.243 0.618 0.347 0.173 0.209 0.154
MSE@60% 0.072 0.086 0.170 0.121 0.135 0.133 0.840 0.210 0.061 0.092 0.055
MAE@70% 0.205 0.218 0.218 0.186 0.325 0.196 0.798 0.413 0.174 0.246 0.172
MSE@70% 0.093 0.104 0.111 0.083 0.228 0.085 1.303 0.303 0.071 0.132 0.069
ETTm1 MAE@0% 0.2265 0.2405 0.2309 0.2379 0.2772 0.2423 0.4759 0.3331 0.2999 0.2100 0.2000
MSE@0% 0.2302 0.2741 0.2514 0.2648 0.3504 0.2782 0.8854 0.4218 0.2682 0.2018 0.2008
MAE@20% 0.2616 0.2900 0.2909 0.2779 0.3711 0.2971 0.5198 0.3809 0.3271 0.2584 0.2368
MSE@20% 0.3068 0.3593 0.3751 0.3486 0.5499 0.3864 0.9345 0.5177 0.3282 0.2992 0.2586
MAE@40% 0.3005 0.3308 0.3593 0.3456 0.4418 0.3592 0.5787 0.4391 0.3571 0.3040 0.2734
MSE@40% 0.4007 0.4645 0.4851 0.4671 0.7412 0.4801 1.0499 0.6643 0.4749 0.4093 0.3382
MAE@60% 0.3678 0.4171 0.4365 0.4091 0.5551 0.4191 0.7038 0.5457 0.4315 0.3993 0.3414
MSE@60% 0.5960 0.7328 0.7580 0.6769 1.1067 0.7022 1.3801 0.9573 0.6978 0.6525 0.5200
MAE@70% 0.4266 0.4863 0.5034 0.4743 0.6508 0.4816 0.8287 0.6278 0.4845 0.4828 0.4013
MSE@70% 0.8006 0.9928 1.0663 0.9325 1.4721 0.9837 1.7733 1.2710 0.9035 0.8988 0.6956
ETTm2 MAE@0% 0.0802 0.0821 0.0781 0.0824 0.0918 0.0828 0.1313 0.1327 0.0813 0.0782 0.0746
MSE@0% 0.0175 0.0182 0.0161 0.0182 0.0217 0.0187 0.0410 0.0363 0.0160 0.0165 0.0152
MAE@20% 0.0956 0.1059 0.1020 0.1005 0.1405 0.1024 0.2687 0.1760 0.0888 0.1087 0.0885
MSE@20% 0.0231 0.0266 0.0284 0.0251 0.0391 0.0278 0.1976 0.0591 0.0222 0.0260 0.0203
MAE@40% 0.1159 0.1355 0.2673 0.2041 0.1743 0.2441 0.4131 0.1984 0.1088 0.1417 0.1037
MSE@40% 0.0305 0.0394 0.1688 0.1210 0.0601 0.1490 0.4003 0.0765 0.0267 0.0406 0.0258
MAE@60% 0.1432 0.1790 0.2651 0.2425 0.2412 0.2398 0.6221 0.2479 0.1360 0.1986 0.1266
MSE@60% 0.0477 0.0685 0.1671 0.1423 0.1215 0.1448 0.8293 0.1194 0.0424 0.0811 0.0384
MAE@70% 0.1650 0.1899 0.1905 0.1803 0.3104 0.1728 0.8024 0.3130 0.1602 0.2317 0.1467
MSE@70% 0.0619 0.0808 0.0861 0.0779 0.2056 0.0717 1.2968 0.1857 0.0569 0.1142 0.0520
Weather MAE@0% 0.031 0.030 0.030 0.031 0.034 0.031 0.051 0.051 0.035 0.028 0.028
MSE@0% 0.021 0.019 0.018 0.019 0.023 0.020 0.042 0.029 0.020 0.016 0.016
MAE@20% 0.042 0.050 0.044 0.048 0.099 0.046 0.173 0.073 0.041 0.062 0.038
MSE@20% 0.028 0.025 0.031 0.027 0.045 0.032 0.220 0.036 0.028 0.033 0.025
MAE@40% 0.050 0.074 0.152 0.150 0.135 0.157 0.300 0.095 0.051 0.088 0.050
MSE@40% 0.037 0.036 0.167 0.156 0.078 0.140 0.500 0.060 0.033 0.047 0.033
MAE@60% 0.066 0.102 0.175 0.169 0.197 0.168 0.466 0.150 0.066 0.128 0.062
MSE@60% 0.061 0.062 0.223 0.169 0.178 0.163 1.099 0.150 0.064 0.089 0.057
MAE@70% 0.076 0.121 0.101 0.118 0.254 0.100 0.586 0.204 0.076 0.166 0.070
MSE@70% 0.081 0.084 0.138 0.128 0.318 0.120 1.707 0.284 0.083 0.154 0.075
Exchange MAE@0% 0.0185 0.0187 0.0190 0.0186 0.0205 0.0187 0.0264 0.3223 0.0344 0.0195 0.0184
MSE@0% 0.0010 0.0010 0.0010 0.0010 0.0011 0.0010 0.0017 0.1724 0.0031 0.0011 0.0009
MAE@20% 0.0292 0.0324 0.0288 0.0273 0.0795 0.0227 0.1992 0.2969 0.0364 0.0458 0.0217
MSE@20% 0.0031 0.0024 0.0024 0.0019 0.0126 0.0016 0.2089 0.1592 0.0029 0.0041 0.0016
MAE@40% 0.0629 0.0783 0.1899 0.1416 0.1469 0.1408 0.4358 0.3326 0.0431 0.0786 0.0253
MSE@40% 0.0129 0.0127 0.0770 0.0884 0.0429 0.1047 0.5718 0.2146 0.0034 0.0130 0.0015
MAE@60% 0.1165 0.1364 0.2439 0.1978 0.2599 0.1283 0.7911 0.4334 0.0608 0.1357 0.0363
MSE@60% 0.0363 0.0405 0.1060 0.1129 0.1394 0.0831 1.3714 0.3814 0.0071 0.0404 0.0031
MAE@70% 0.1418 0.1836 0.2865 0.1321 0.3778 0.0718 1.0766 0.5412 0.0846 0.1991 0.0517
MSE@70% 0.0490 0.0697 0.1413 0.0581 0.3011 0.0169 2.2850 0.5912 0.0128 0.0881 0.0058
BeijingAir MAE@0% 0.2552 0.2571 0.2606 0.2609 0.2707 0.2601 0.3431 0.3167 0.2533 0.2583 0.2526
MSE@0% 0.3026 0.3085 0.3156 0.3008 0.3263 0.3102 0.5141 0.3779 0.3189 0.2939 0.2929
MAE@20% 0.2849 0.2896 0.2923 0.2866 0.3204 0.2949 0.4035 0.3478 0.2770 0.2922 0.2758
MSE@20% 0.3467 0.3651 0.3655 0.3421 0.3853 0.3603 0.6010 0.4196 0.3496 0.3488 0.3317
MAE@40% 0.3188 0.3312 0.3545 0.3278 0.3620 0.3480 0.4714 0.3950 0.3297 0.3322 0.3106
MSE@40% 0.4285 0.4601 0.4820 0.4305 0.4760 0.4689 0.7584 0.5474 0.4645 0.4422 0.4120
MAE@60% 0.3860 0.4011 0.4124 0.3976 0.4362 0.4142 0.5930 0.4750 0.4006 0.4017 0.3728
MSE@60% 0.5939 0.6503 0.6573 0.6458 0.6799 0.6624 1.1378 0.7803 0.6110 0.6275 0.5862
MAE@70% 0.4412 0.4632 0.4691 0.4613 0.5022 0.4673 0.7057 0.5456 0.4549 0.4789 0.4337
MSE@70% 0.7595 0.8730 0.8874 0.8540 0.9358 0.8745 1.5761 1.0151 0.8327 0.9234 0.8010